# Conformal Generative Modeling with Improved Sample Efficiency through Sequential Greedy Filtering

### v1

@@ 1 Introduction
Generative models have found extensive applications across various domains, including the generation of images (Rombach et al., 2022), molecules (Vignac et al., 2023), and natural language (Achiam et al., 2023). While these models are becoming increasingly performant, caution is required when decisions are based on generated outputs. This is particularly relevant for scenarios where errors have severe consequences. A prime example is the phenomenon of “confabulations” (or “hallucintations”) in large language models—responses that lack factual support, potentially leading to substantial harm if trusted (Bai et al., 2024; Weidinger et al., 2021). Therefore, developing effective measures to ensure strong statistical guarantees on the predictions of these models is imperative.
@@
Figure 1: SCOPE-Gen for Radiology Report Generation. The sequential prediction procedure can be separated into two stages: In the generation stage ($s=0$), i.i.d. text reports (blue tokens) are drawn from the generative model. In the filter stage $(s\in\{1,2\})$, the prediction set from the previous step $s-1$ is refined to remove examples with low quality (removing the false response in red) and examples with high similarity to already sampled texts (the answers in green are similar). Each step is performed via iterative (sub-)sampling in a given order (indicated by the purple circles), until the value of a non-conformity threshold $\lambda_{(s)}^{*}$ is exceeded.
Addressing this problem, Quach et al. (2024) have recently introduced a method called conformal language modeling (CLM) [^1] Notably, despite its name, CLM is not based on conformal prediction in a strict sense.. Rather than evaluating a non-conformity measure on each point in the prediction space (which is intractable for textual data), CLM calibrates a stopping rule parameter $\lambda_{(0)}$ that decides how many samples $y$ are required from a generative model $G$, conditionally on features $x$ (for example, an image or a text query). These samples then make up the prediction set, which fulfills a statistical guarantee with respect to containing at least one admissible (correct/non-hallucinated) example. The method further introduces rejection sampling to ensure that the prediction set satisfies desired properties. The sampling is controlled by two calibration parameters $\lambda_{(1)}$ and $\lambda_{(2)}$ that ensure
@@
While CLM offers a notion of admissibility control in generative models, it has a practical limitation. Specifically, during calibration, CLM necessitates sampling and evaluating admissibility for a fixed number of times per instance $x$ in the calibration set. In safety-critical applications, such as medicine, the factual integrity of admissibility checks based on n-gram (Lin, 2004; Papineni et al., 2002) or machine learning methods (Smit et al., 2020) remains a concern (Zhang et al., 2024), necessitating manual assessments by human domain experts for accurate calibration. Reducing the number of these costly admissibility checks is thus paramount (Settles, 2009; Mosqueira-Rey et al., 2023).
@@
Section § 2 introduces (conformal) admissibility control, which is the main objective of the present work. Thereafter, we present conformal prediction in Section § 3. These sections set the stage for presenting our method in Section § 4. Upon establishing the main principles, we demonstrate the sub-sampling and non-conformity update functions used for filtering (§ 4.1), the algorithmic details of the calibration (§ 4.2). Thereafter, we experimentally demonstrate the advantages of our approach in comparison to prior work in Section § 5, followed by a section dedicated to related work (§ 6). Finally, we discuss limitations in Section § 7 and conclude in Section § 8.
@@ 2 Objective
where $(X_{n+1},Y^{t}_{n+1})$ is an additional independent sample from the same distribution as the calibration set and $a:\mathcal{Y}^{2}\mapsto\{0,1\}$ measures admissibility between generated sample $Y\in\mathcal{Y}$ and $Y^{t}_{n+1}\in\mathcal{Y}$ (Fisch et al., 2021). For instance, $a$ could measure whether a textual response from a language model $Y$ for the input $X$ is similar to a ground truth textual response $Y_{n+1}^{t}$. For completeness, we demonstrate the relationship between admissibility control 1 and the more general notion called conformal risk control (Angelopoulos et al., 2022) in App. A. For notational convenience, we introduce the set-level admissibility
@@ 3 Preliminary : Conformal Prediction
where $\widehat{\text{quantile}}$ denotes the empirical quantile of $\nu$, evaluated at the calibration points $\left\{\nu(x_{i},y_{i})\right\}_{i=1}^{n}$.
@@ 4 Sequential Conformal Prediction for Generative Models
SCOPE-Gen relies on i.i.d. calibration data $\{(x_{i},y^{t}_{i})\}_{i=1}^{n}$ from the true generating process. In order to control $\mathcal{A}(\bm{\lambda})$ 3 at level $\alpha$, we utilize conformal prediction § 3. The main challenge of our setting is that the output space $\mathcal{Y}$ is combinatorially large or even infinite. This means that the prediction set determined as in 5 (with $\bm{\lambda}=\lambda\in\mathbb{R}$) is intractable. Thus, we construct a prediction set directly from i.i.d. samples from a generative model $G$.
@@
    Generation Stage ($s=0$): An initial prediction set $\mathcal{C}_{(0)}$ is generated by iteratively adding new i.i.d. samples from a generative model $G$, conditioned on $x$ (Alg. 1 and Alg. 1). A new sample $y$ is added until the value of a non-conformity measure $\nu$, which is updated at each iteration of the while loop (Alg. 1), exceeds the value of the calibration parameter $\lambda_{(0)}$ (Alg. 1). We assume that $\nu$ is increasing at each iteration.
@@ 18             $j\leftarrow j+1$
This sequential prediction provides a remedy for the intractibility of the conformal prediction set as computed in 5. However, it poses the challenge that the total admissibility $\mathcal{A}(\bm{\lambda})$ is parameterized by a multi-dimensional calibration parameter $\bm{\lambda}$, which cannot be directly calibrated via conformal prediction 6. This is the case because conformal prediction crucially depends on the assumption that the loss is controlled by a one-dimensional parameter (Angelopoulos et al., 2022).
@@
This means that the probability that the final prediction set $\mathcal{C}_{(2)}$ contains an admissible example is determined by the probability that the first step generates at least one admissible example and the probabilities that none of the consecutive steps filter out all admissible examples. Consequently, we can control the total admissibility $\mathcal{A}(\bm{\lambda})$ at level $\alpha$ by controlling each individual admissibility $\mathcal{A}(\lambda_{(s)})$ for univariate $\lambda_{(s)}\in\mathbb{R}$ and for $s\in\{0,1,2\}$, separately, where the individual levels $\alpha_{(s)}$ are chosen such that $\alpha=1-\prod_{s=0}^{2}\left(1-\alpha_{(s)}\right)$. We discuss the choice of $\alpha_{(s)}$ in § C.1.
@@
where $\lambda^{\ast}$ is the calibrated parameter 6 and we recall that $\nu$ is increasing in $l$, for fixed $x$ [^3] We note that computing $\tilde{\mathcal{C}}_{(s)}(x)$ in 9 would generally be intractable for non-increasing $\nu$.. We see that sampling the prediction set $\mathcal{C}_{(s)}$ by sampling $\mathop{\mathrm{argmax}}_{l}\tilde{\mathcal{C}}_{(s)}$ many examples controls the individual admissibility $\mathcal{A}(\lambda_{(s)})$ at level $\alpha_{(s)}$, which is equivalent to adding new samples to $\mathcal{C}_{(s)}$ until the non-conformity measure $\nu$ exceeds $\lambda_{(s)}$, as done in Alg. 1.
@@ 4.1 Greedy Filters
where $d_{\text{max}}$ is an upper bound on the distance function (see § C.2 for details).
@@
Using a quality function $q$, we choose the point $\mathcal{C}_{(s-1)}\backslash\mathcal{C}_{(s)}$ with the maximum quality:
@@ 4.2 Calibration
As outlined (§ 4), for each $x_{i}$, we generate i.i.d. samples $y_{i}^{(j)}$ from $G\,|\,x_{i}$, until the first $y_{i}^{(j)}$ is admissible (Alg. 2). The resulting non-conformity values $\nu_{i}$ are then used to calibrate $\lambda_{(0)}$. In practice, we typically need to limit the maximum amount of samples drawn by $\texttt{max}\in\mathbb{N}$ (Alg. 2) to tackle “hard” cases, i.e., examples $(x_{i},y_{i}^{t})$ where the generative model is unable to generate an admissible prediction. For these examples, we cannot generate finite non-conformity values and set these values to $\infty$ (Alg. 2). Provided that such examples occur infrequently, the conformal quantile $\lambda^{\ast}_{(0)}$ 6 is not affected. Otherwise, SCOPE-Gen rejects the calibration444which amounts to returning the entire output space $\mathcal{Y}$., as common for methods in the field of risk control (e.g., Angelopoulos et al. (2021)).
@@ 5 Experiments
We compare our method against prior work and derived baselines to empirically demonstrate the advantages of SCOPE-Gen on real-world natural language tasks and a molecular generation task. As the motivation of the present work lies in sample efficiency, we generally focus on small calibration sets. The experiments in the main text consider the sum non-conformity measure with $\gamma=0.5$ 10. We show qualitative evaluations in App. F, further quantitative evaluations in § E.2 and an analysis of empirically achieved admissibility 3 in § E.1.
We compare our method against conformal language modeling (CLM; Quach et al. (2024)) and a baseline derived from it. CLM controls a different notion of admissibility than we do. Specifically, CLM controls
@@
where $D_{\text{cal}}$ is shorthand for the calibration data $\{(x_{i},y_{i}^{t})\}_{i=1}^{n}$. We note that 12 controls admissibility 1 at level $\alpha$ if $\alpha=\beta_{(1)}+\beta_{(2)}-\beta_{(1)}\beta_{(2)}$ (we show this in App. B). In order to tighten the gap, we perform a grid search over combinations of $\beta_{(1)}$ and $\beta_{(2)}$ that correspond to $\alpha$ and choose the configuration that minimizes admissibility (more details in § C.3). We do this for each experiment separately and compare the optimal configuration to our method at level $\alpha$.
@@
Baselines.
1) CLM: Conformal Language Modeling (Quach et al., 2024). 2) CLM reduced max: This baseline assesses how CLM performs if we enforce a lower amount of admissibility checks during calibration by limiting max. Specifically, we choose $\texttt{max}=10$ as limit. 3) SCOPE-Gen gen.-only: We only perform SCOPE-Gen using the generation stage and do not apply any filters. 4) SCOPE-Gen flipped: We apply the quality filter first and the diversity filter second (§ 4.1; assessed when both filters are used). We present baselines 3) and 4) in the additional experiments (App. E).
@@ 5.1 Setup
For sake of demonstration, all experiments shown here use automated admission functions. We leave experiments that involve admissibility checks made by human domain experts for future work. We only provide a general description in the main text and refer to § D.1 for details regarding the setup and § D.2 for illustrative examples of each task.
@@ 5.2 Results
SCOPE-Gen consistently outperforms CLM in terms of amount of admissibility checks (# Queries) and runtime (Time), as can be seen in Tab. 1. For instance, on MIMIC-CXR, SCOPE-Gen requires only $\frac{4.21}{20}\approx 21\%$ of the admissibility queries required for CLM. The “CLM reduced max” baseline demonstrates that counteracting this disadvantage by reducing the maximum sample size max comes at a high price to pay: The amount of rejected calibrations increases greatly (e.g., $81.7\%$ on MIMIC-CXR), while still being far from SCOPE-Gen in terms of performance. This decreased performance is due to the fact that the admissibility guarantee in CLM holds for this specified max parameter, also during inference. In the inference phase, however, an upper bound on the prediction set size is often irrelevant, because no more admissibility checks must be performed. In addition, SCOPE-Gen consistently generates smaller prediction sets when the count or sum non-conformity updates are chosen (see 10). The only setting where CLM outperforms SCOPE-Gen in terms of prediction set size is for the max baseline, but not across all experiments (shown in App. E).
Table 1: Quantitative Evaluations. All metrics $\pm$ standard deviation of the different baselines from $300$ repeated evaluations. For each evaluation, we first sub-sample a calibration set of size $n=600$ and calibrate each method at coverage level $\alpha=0.3$ with sum non-conformity 10. Then, we take means for each metric on a test set of size $300$, sampled from the remaining data. Best is bold.
@@ 6 Related Work
Works other than the one by Quach et al. (2024) that target language generation with a different scope to ours include the approach by Ulmer et al. (2024) who consider language modeling for non-exchangeable data and  Gui et al. (2024) who focus on conformal alignment of language models.
7 Discussion & Limitations
Data Splitting.
Conformal Prediction vs. Learn-then-Test.

### v2

@@ 1 Introduction
Generative models have found extensive applications across various domains, including the generation of images (Rombach et al., 2022), molecules (Vignac et al., 2023), and natural language (Achiam et al., 2023). While these models are becoming increasingly performant, caution is required when decisions are based on generated outputs. This is particularly relevant for scenarios where errors have severe consequences. A prime example is the phenomenon of “confabulations” (or “hallucinations”) in large language models—responses that lack factual support, potentially leading to substantial harm if trusted (Bai et al., 2024; Weidinger et al., 2021). Therefore, developing effective measures to ensure strong statistical guarantees on the predictions of these models is imperative.
@@
Figure 1: SCOPE-Gen for Radiology Report Generation. The sequential prediction procedure can be separated into two stages: In the generation stage ($s=0$), i.i.d. text reports (blue tokens) are drawn from the generative model. In the filter stage $(s\in\{1,2\})$, the prediction set from the previous step $s-1$ is refined to remove examples with low quality (removing the false response in red) and examples with high similarity to already sampled texts (the answers in green are similar). Each step is performed via iterative (sub-)sampling in a given order (indicated by the purple circles), until the value of a non-conformity variable $\nu$ exceeds a calibrated threshold $\lambda_{(s)}^{*}$.
Quach et al. (2024) have recently introduced a method called conformal language modeling (CLM) for addressing this problem.  [^1] Notably, despite its name, CLM is not based on conformal prediction in a strict sense. Rather than evaluating a non-conformity measure on each point in the prediction space (which is intractable for textual data), CLM calibrates a stopping rule parameter $\lambda_{(0)}$ that decides how many samples $y$ are required from a generative model $G$, conditionally on features $x$ (for example, an image or a text query). These samples then make up the prediction set, which fulfills a statistical guarantee with respect to containing at least one admissible (correct/non-hallucinated) example. The method further introduces rejection sampling to ensure that the prediction set satisfies desired properties. The sampling is controlled by two calibration parameters $\lambda_{(1)}$ and $\lambda_{(2)}$ that ensure
@@
While CLM offers a notion of admissibility control in generative models, it has a practical limitation. Specifically, during calibration, CLM necessitates sampling and evaluating admissibility for a fixed number of times per instance $x$ in the calibration set. In safety-critical applications, such as medicine, the factual integrity of admissibility checks based on n-gram (Lin, 2004; Papineni et al., 2002) or machine learning methods (Smit et al., 2020) remains a concern (Zhang et al., 2024b), necessitating manual assessments by human domain experts for accurate calibration. Reducing the number of these costly admissibility checks is thus paramount (Settles, 2009; Mosqueira-Rey et al., 2023).
@@
Section § 2 introduces (conformal) admissibility control, which is the main objective of the present work. Thereafter, we present conformal prediction in Section § 3. These sections set the stage for presenting our method in Section § 4. Upon establishing the main principles, we demonstrate the sub-sampling and non-conformity update functions used for filtering (§ 4.1), the algorithmic details of the calibration (§ 4.2). Thereafter, we experimentally demonstrate the advantages of our approach in comparison to prior work in Section § 5, followed by a section dedicated to related work (§ 6). Finally, we conclude in Section § 7. A discussion section is provided in App. I.
@@ 2 Objective
where $(X_{n+1},Y^{t}_{n+1})$ is an additional independent sample from the same distribution as the calibration set and $a:\mathcal{Y}^{2}\mapsto\{0,1\}$ (no additional assumptions on $a$ are imposed) measures admissibility between generated sample $Y\in\mathcal{Y}$ and $Y^{t}_{n+1}\in\mathcal{Y}$ (Fisch et al., 2021)333 We note that $Y^{t}_{n+1}$ is not observed at test time. . For instance, $a$ could measure whether a textual response from a language model $Y$ for the input $X$ is similar to a ground truth textual response $Y_{n+1}^{t}$. For completeness, we demonstrate the relationship between admissibility control 1 and the more general notion called conformal risk control (Angelopoulos et al., 2022) in App. A. For notational convenience, we introduce the set-level admissibility
@@ 3 Preliminary : Conformal Prediction
where $\widehat{\text{quantile}}$ denotes the empirical quantile of $\nu$, evaluated at the calibration points $\left\{\nu(x_{i},y_{i})\right\}_{i=1}^{n}$444 Intuitively, $\lambda^{*}$ is the quantile at level $\approx 1-\alpha$, with a small correction for theoretical reasons (e.g., Angelopoulos & Bates (2021)). . We provide a rigorous definition of the empirical quantile 6 in App. B .
@@ 4 Sequential Conformal Prediction for Generative Models
SCOPE-Gen relies on i.i.d. calibration data $\{(x_{i},y^{t}_{i})\}_{i=1}^{n}$ from the true generating process. In order to control $\mathcal{A}(\bm{\lambda})$ 3 at level $\alpha$, we utilize conformal prediction § 3. The main challenge of our setting is that the output space $\mathcal{Y}$ is combinatorially large or even infinite. Hence, the prediction set determined as in 5 (with $\bm{\lambda}=\lambda\in\mathbb{R}$) is intractable. Thus, we construct a prediction set directly from i.i.d. samples from a generative model $G$.
@@
    Generation Stage ($s=0$): An initial prediction set $\mathcal{C}_{(0)}$ is generated by iteratively adding new i.i.d. samples from a generative model $G$, conditioned on $x$ (Alg. 1, Alg. 1). A new sample $y$ is added until a non-conformity value $\nu$, which is updated at each iteration of the while loop (Alg. 1), exceeds the value of the calibration parameter $\lambda_{(0)}$ (Alg. 1). We assume that $\nu$ is increasing at each iteration.
@@ 18             $j\leftarrow j+1$
This sequential prediction provides a remedy for the intractibility of the conformal prediction set as computed in 5. However, it poses the challenge that the total admissibility $\mathcal{A}(\bm{\lambda})$ is parameterized by a multi-dimensional calibration parameter $\bm{\lambda}$, which cannot be directly calibrated via conformal prediction 6. This is due to the fact that conformal prediction crucially depends on the assumption that the loss is controlled by a one-dimensional parameter (Angelopoulos et al., 2022).
@@
This means that the probability that the final prediction set $\mathcal{C}_{(2)}$ contains an admissible example is determined by the probability that the first step generates at least one admissible example and the probabilities that none of the consecutive steps filter out all admissible examples. Consequently, we can control the total admissibility $\mathcal{A}(\bm{\lambda})$ at level $\alpha$ by controlling each individual admissibility $\mathcal{A}(\lambda_{(s)})$ for univariate $\lambda_{(s)}\in\mathbb{R}$ and for $s\in\{0,1,2\}$, separately, where the individual levels $\alpha_{(s)}$ are chosen such that $\alpha=1-\prod_{s=0}^{2}\left(1-\alpha_{(s)}\right)$. We discuss the choice of $\alpha_{(s)}$ in § D.1.
@@
where $\lambda^{\ast}$ is the calibrated parameter 6 and we recall that $\nu$ is increasing in $l$, for fixed $x$ [^5] We note that computing $\tilde{\mathcal{C}}_{(s)}(x)$ in 9 would generally be intractable for non-increasing $\nu$. We see that sampling the prediction set $\mathcal{C}_{(s)}$ by sampling $\mathop{\mathrm{argmax}}_{l}\tilde{\mathcal{C}}_{(s)}$ many examples controls the individual admissibility $\mathcal{A}(\lambda_{(s)})$ at level $\alpha_{(s)}$, which is equivalent to adding new samples to $\mathcal{C}_{(s)}$ until the non-conformity measure $\nu$ exceeds $\lambda_{(s)}$, as done in Alg. 1.
@@ 4.1 Greedy Filters
where $d_{\text{max}}$ is an upper bound on the distance function (see § D.2 for details).
@@
Using a quality function $q$, we choose the point $\mathcal{C}_{(s-1)}\backslash\mathcal{C}_{(s)}$ with maximum quality:
@@ 4.2 Calibration
As outlined (§ 4), for each $x_{i}$, we generate i.i.d. samples $y_{i}^{(j)}$ from $G\,|\,x_{i}$, until the first $y_{i}^{(j)}$ is admissible (Alg. 2). The resulting non-conformity values $\nu_{i}$ are then used to calibrate $\lambda_{(0)}$. In practice, we typically need to limit the maximum amount of samples drawn by $\texttt{max}\in\mathbb{N}$ (Alg. 2) to tackle “hard” cases, i.e., examples $(x_{i},y_{i}^{t})$ where the generative model is unable to generate an admissible prediction. For these examples, we cannot generate finite non-conformity values and set these values to $\infty$ (Alg. 2). Provided that such examples occur infrequently, the conformal quantile $\lambda^{\ast}_{(0)}$ 6 is not affected. Otherwise, SCOPE-Gen rejects the calibration666which amounts to returning the entire output space $\mathcal{Y}$., as common for methods in the field of risk control (e.g., Angelopoulos et al. (2021)).
@@ 5 Experiments
We compare our method against prior work and derived baselines to empirically demonstrate the advantages of SCOPE-Gen on real-world natural language tasks and a molecular generation task. As the motivation of the present work lies in sample efficiency, we generally focus on small calibration sets. The experiments in the main text consider the sum non-conformity measure with $\gamma=0.5$ 10. We show qualitative evaluations in App. G, further quantitative evaluations and ablation studies in § F.2 and an analysis of empirical admissibility 3 in § F.1.
We compare our method against CLM (Quach et al., 2024) and a baseline derived from it. CLM controls a different notion of admissibility than we do. Specifically, CLM controls
@@
where $D_{\text{cal}}$ is shorthand for the calibration data $\{(x_{i},y_{i}^{t})\}_{i=1}^{n}$. We note that 12 controls admissibility 1 at level $\alpha$ if $\alpha=\beta_{(1)}+\beta_{(2)}-\beta_{(1)}\beta_{(2)}$ (we show this in App. C). In order to tighten the gap, we perform a grid search over combinations of $\beta_{(1)}$ and $\beta_{(2)}$ that correspond to $\alpha$ and choose the configuration that minimizes admissibility (more details in § D.3). We do this for each experiment separately and compare the optimal configuration to our method at level $\alpha$.
@@
Baselines & Ablations .
1) CLM (Quach et al., 2024). 2) CLM reduced max: This baseline assesses how CLM performs if we enforce a lower amount of admissibility checks during calibration by limiting max. Specifically, we choose $\texttt{max}=10$ as limit. 3) SCOPE-Gen gen.-only: We only perform SCOPE-Gen using the generation stage and do not apply any filters. 4) SCOPE-Gen flipped: We apply the quality filter first and the diversity filter second (§ 4.1; config. 1 [§ D.1]; when both filters are used). We present 3) and 4) in App. F.
@@ 5.1 Setup
For sake of demonstration, experiments shown here use automated admission functions. We leave experiments that involve admissibility checks made by human domain experts for future work777 We stress that admissibility control 1 still applies if a human domain expert assesses admissibility. . We provide a general description in the main text and refer to § E.1 regarding setup and § E.2 for examples of each task. For SCOPE-Gen, we show config. 1 (see § D.1) in the main text.
@@ 5.2 Results
SCOPE-Gen consistently outperforms CLM in terms of amount of admissibility checks (# Queries) and runtime (Time), as can be seen in Tab. 1. For instance, on MIMIC-CXR, SCOPE-Gen requires only $\frac{4.21}{20}\approx 21\%$ of the admissibility queries required for CLM. The “CLM reduced max” baseline demonstrates that counteracting this disadvantage by reducing the maximum sample size max comes at a high price to pay: The amount of rejected calibrations increases greatly (e.g., $81.7\%$ on MIMIC-CXR), while still being far from SCOPE-Gen in terms of performance. This decreased performance is due to the fact that the admissibility guarantee in CLM holds for this specified max parameter, also during inference. In the inference phase, however, an upper bound on the prediction set size is often irrelevant, because no more admissibility checks must be performed. In addition, SCOPE-Gen consistently generates smaller prediction sets when the count or sum non-conformity updates are chosen (see 10). The only setting where CLM outperforms SCOPE-Gen in terms of prediction set size is for the max baseline, but not across all experiments (shown in App. F). An additional benefit of SCOPE-Gen in comparison to CLM is the vast reduction in computational demand for calibration: For MIMIC-CXR, SCOPE-Gen takes on average only $0.072$ seconds for a single calibration, in comparison to CLM which takes $113.677$ seconds on average ($\frac{0.072}{113.677}\approx 10^{-4}\cdot 6.3$).
Table 1: Quantitative Evaluations. All metrics ±$\pm 1$ standard deviation of the different baselines from $300$ repeated evaluations. For each evaluation, we first sub-sample a calibration set of size $n=600$ and calibrate each method at coverage level $\alpha=0.3$ with sum non-conformity 10. Then, we take means for each metric on a test set of size $300$, sampled from the remaining data. Best is bold.
@@ 6 Related Work
Works other than the one by Quach et al. (2024) that target language generation with a different scope to ours include the approach by Ulmer et al. (2024) who consider language modeling for non-exchangeable data, Gui et al. (2024) who focus on conformal alignment of language models and Kumar et al. (2023) who specifically consider multiple choice problems. Another concurrent line of work, followed by Cherian et al. (2024); Mohri & Hashimoto (2024); Liu & Wu (2024), provides guarantees on individual long-form text outputs by chunking them into sub-claims and removing false sub-claims. A critical assumption for such methods to be practical is that text responses indeed consist of multiple distinct claims with little coherence.
Detecting Hallucination without Guarantees.
There is a body of work focused on quantifying uncertainty and detecting hallucinations in language models, though these approaches do not directly achieve distribution-free risk control. Existing methods in this domain are based on notions such as semantic (Kuhn et al., 2023; Lin et al., 2024) or task-specific uncertainty (Wang & Holmes, 2024) and self-consistency (Zhang et al., 2024a). Although not explored in the present work, we believe that such approaches offer a promising opportunity for cross-pollination with ideas from conformal prediction in future work.